OpenStudy (anonymous):

okay not completely sure what I need to do. All I have is find x and y intercepts. Using calculus

OpenStudy (anonymous):

\(x\) intercepts is when ALL OTHER variables are set to 0, and you solve for \(x\). Same thing applies to \(y\).

OpenStudy (anonymous):

Could you do something like slope formula and once you have the slope plug in to the y=mx+b formula?

OpenStudy (anonymous):

that would be if I was given the lines only

OpenStudy (anonymous):

What is the equation? I need context.

OpenStudy (anonymous):

wasn't given an equation just what I need to know how to do

OpenStudy (anonymous):

Then come up with a practice question on your own.

OpenStudy (anonymous):

given the line x-3 and -x-3 find the x and y intercepts

OpenStudy (anonymous):

as @wio said, the intercepts occur when the other variables are 0. if y = x - 3 the x intercept is when y = 0 and the y intercept is when x = 0